OpenStudy (antiprime):

H(x)=(x^2)+1 and K(x)=-(x^2)+4. If K(H)=0, what are the roots/solutions? a. ±i√3 b. ±i c. ±2 d. ±1 e. ±1, ±i√3

OpenStudy (antiprime):

help

OpenStudy (mayankdevnani):

have you tried it ?

OpenStudy (antiprime):

i get c

OpenStudy (mayankdevnani):

can you show your work ?

OpenStudy (antiprime):

ill try

OpenStudy (misty1212):

hI!!

OpenStudy (misty1212):

ok that was totally wrong you have to find \[K(H(x))\]first

OpenStudy (antiprime):

h(x)=X^2+1; k(x)=-x^2=4 k(h)=k(x^2+1)=-(x^2+1)^2+4=0 (x^2+1)^2 x^2 +1= pm 2 x^2+1=-2 or x^2+1=2 \[x=\pm i \sqrt{3}\]

OpenStudy (antiprime):

that????

OpenStudy (mayankdevnani):